OpenStudy (anonymous):

Quadrilateral ABCD has coordinates A (3, -5), B (5, -2), C (10, -4), D (8, -7). Quadrilateral ABCD is a

OpenStudy (anonymous):

Check the angles to see if they are 90 degrees. Opposite sides in parallelogram share the same length as well.

OpenStudy (anonymous):

AB = sqr root (5-3)^2 + (-2+5)^2 = sqr root 13 BC = sqr root(10-5)^2 + (-4+2)^2= sqr root 29 CD= sqr root(8-10)^2 + ((-7+4)^2 = sqr root 13 DA = sqr root(3-8)^2 + (-5+7)^2 = sqr root 29 ABCD is paralellogram

OpenStudy (mathstudent55):

Now that you know it's a parallelogram, you need to find out if it is also a rectangle. Find the slopes of two adjacent sides. If their product is -1, the sides are perpendicular, and it is a rectangle, if not, then it's just a parallelogram.